Find the new equation of straight line 2x+y=5 after rotating the axes through 45°.
1
Expert's answer
2020-11-29T16:33:35-0500

We have to rotate the axes to "45^0". Let say in anti-clockwise manner.

So,

"x=x\cos \theta-y \sin \theta" , "y=x \sin \theta+y \cos \theta" .


"\implies x=x \cos 45^0-y \sin 45^0=\frac{\sqrt2}{2}x - \frac{\sqrt2}{2}y"

Also,

"y=x\sin 45^0 + y \cos 45^0= \frac{\sqrt2}{2}x + \frac{\sqrt2}{2}y"


Put the new values of "x" and "y" into the straight line equation.


"\implies 2(\frac{\sqrt2}{2}x - \frac{\sqrt2}{2}y)+(\frac{\sqrt2}{2}x + \frac{\sqrt2}{2}y)=5\\
\implies x\sqrt2-y\sqrt2+\frac{\sqrt2}{2}x + \frac{\sqrt2}{2}y=5\\
\implies \frac{3\sqrt2}{2}x - \frac{\sqrt2}{2}y=5"


The new equation of the line is "\frac{3\sqrt2}{2}x - \frac{\sqrt2}{2}y=5."